# Vanishing conductivity-like Gilbert damping in mirror-symmetric van der Waals ferromagnets

### v1

@@ Abstract
The identification of two-dimensional van der Waals ferromagnetic materials has significantly expanded the realm of magnetic materials and enabled innovative control techniques such as gating and stacking. The dynamical behavior of magnetization is profoundly influenced by the Gilbert damping parameter, a crucial factor for enhancing the speed of magnetic switching and reducing the energy consumption of magnetic devices. Despite its importance, the understanding of Gilbert damping in van der Waals ferromagnets remains limited, impeding their technological applications. Here we present a theoretical calculation of Gilbert damping in two-dimensional van der Waals metals, focusing on Fe3GaTe2 and Fe3GeTe2 as model systems. We discover that mirror symmetry prohibits intraband contributions, resulting in disappearance of conductivity-like damping. Consequently, at low temperatures, the Gilbert damping in single-layer Fe3GaTe2 and Fe3GeTe2 is remarkably low when the magnetization is perpendicular to the atomic layer, but it increases substantially when the magnetization is reoriented into the atomic plane. Furthermore, topological nodal lines, also protected by mirror symmetry, contribute significantly to damping mediated by interband transitions and can be effectively modulated by adjusting the Fermi level. Our findings elucidate the distinctive characteristics of Gilbert damping in two-dimensional van der Waals ferromagnets, providing valuable insights for the design and optimization of low-dimensional spintronic devices.
@@
Introduction.—The discovery of two-dimensional (2D) van der Waals (vdW) ferromagnetic (FM) materials, exemplified by CrI3 [1], Cr2Ge2Te6 [2], and Fe3GeTe2 [3, 4], has not only dramatically broadened the horizon of FM materials but also ignited promising prospects for low-dimensional spintronics devices [5, 6]. This promise stems partly from the fact that magnetism in 2D vdW FM metals can be easily controlled through various flexible approaches [7, 8]. As a typical 2D FM metal, Fe3GeTe2 has attracted considerable attention for its spin-orbit torque-driven magnetization switching [9, 10, 11], its unconventional tunability in magnetic tunnel junctions [12, 13] and its potential in novel electronic applications [14]. A very similar vdW FM metal Fe3GaTe2 [15] has been discovered with an almost identical lattice structure to Fe3GeTe2, yet it boasts a Curie temperature above room temperature, rendering it an ideal candidate for integration into magnetic heterostructures and devices that operate under ambient conditions [16, 17, 18, 19].
@@
The quest for designing high-performance memory and logic devices based on 2D vdW FM materials necessitates a profound understanding of the dynamical properties of 2D magnetism [20], particularly how dimensional reduction affects magnetization dynamics in these materials. Gilbert damping, a pivotal parameter in magnetization dynamics [21], dictates the critical current density for current-induced magnetization switching [22], the switching time scale [23] and the velocity of magnetic texture driven by magnetic fields and/or electric currents [24, 25]. In FM transition metals and alloys, such as Fe, Co, Ni [26, 27, 28], FePd [29, 30] and CoFeB [31, 32, 33], the experimentally measured damping exhibits non-monotonic temperature dependence. This behavior is well-understood as the superposition of a conductivity-like component that decreases with temperature and a resistivity-like contribution increasing with temperature. These two components dominate in the low- and high-temperature regimes, respectively. However, in the recently discovered 2D FM metals, Fe3GeTe2 and Fe3GaTe2, only the resistivity-like damping has been observed in the limited experimental [34] and theoretical studies [35]. Furthermore, significant damping anisotropy has also been identified in these 2D FM metals [35, 36], suggesting an atypical magnetization relaxation likely stemming from their low-dimensional lattice structures.
@@
In this article, we present a theoretical calculation of the Gilbert damping in 2D vdW FM metals, with a focus on Fe3GaTe2 and Fe3GeTe2, based on the first-principles electronic structures. We reveal that the mirror symmetry in these vdW systems precludes the intraband transitions, leading to the absence of conductivity-like damping. Furthermore, the symmetry-protected nodal lines near the Fermi level substantially enhance interband transitions, thereby amplifying the resistivity-like damping. Our findings predict a pronounced anisotropy in Gilbert damping when mirror symmetry is broken by rotating the magnetization from an out-of-plane to an in-plane orientation.
Figure 1: (a) Schematic representation of the atomic structure of 1QL Fe3GaTe2. The magnetic moments on the Fe atoms are indicated by black arrows. The blue plane in the middle of the layer highlights the mirror symmetry of the lattice structure. (b) Gilbert damping as a function of the electronic scattering rate for 1QL Fe3GaTe2 with the out-of-plane (red solid circles) and in-plane magnetization (black solid squares). The dotted and dashed lines represent the contributions from intraband and interband transitions to the damping, respectively. The orange and green vertical lines denote two characteristic scattering rates, $\gamma=3$ meV and 20 meV, which are used to calculate the damping in detail.
@@
Theoretical methods.—We utilize the torque-correlation model, as established in literature [37, 38, 39], to calculate the Gilbert damping of 2D vdW FM metals based upon first-principles electronic structure. The damping parameter, expressed as a dimensionless quantity, is given by
[…]
@@
Here $g$, $\mu_{\textrm{B}}$ and $M_{s}$ are the Landé factor, the Bohr magneton and the saturation magnetization, respectively. While indices $n$ and $m$ run over all energy bands, the energy derivative of the Fermi-Dirac distribution $\eta(\epsilon)\equiv-\partial f/\partial\epsilon$ highlights that the damping is predominantly influenced by electronic states in the vicinity of the Fermi surface. The matrix element $\Gamma^{-}_{mn}(\mathbf{k})=\langle\Psi_{m\mathbf{k}}|[\sigma^{-},H_{\textrm{soc}}]|\Psi_{n\mathbf{k}}\rangle$ characterizes the torque on electron spin stemming from spin-orbit coupling (SOC). The spectral function $A_{n\mathbf{k}}(\epsilon)$ is modeled as a Lorentzian centered at the band energy $\epsilon_{nk}$, with its width determined by the electronic scattering rate $\gamma$ [38, 39]. Prior studies have shown that intraband ($n=m$) transitions yield a conductivity-like damping component, which is proportional to the electronic relaxation time, $\tau=\hbar/2\gamma$ [37, 38], while interband ($n\neq m$) transitions contribute to a resistivity-like damping component that scales monotonically with increasing $\gamma$ [40]. To explicitly evaluate these intraband and interband transitions in Eq. (1), we have developed a Wannier interpolation method to perform the integral of $\Gamma^{-}_{mn}(\mathbf{k})$ based on the first-principles electronic structures. The self-consistent band structures are calculated using Quantum ESPRESSO [41], and the Wannier functions are constructed using WANNIER90 [42].
@@
Intraband transitions.—The lattice structure of a single layer Fe3GaTe2, referred to as one quintuple-layer (1QL), is depicted in Fig. 1(a). This lattice, belonging to the D3h point group, maintains mirror symmetry about the blue plane only when the magnetization is perpendicular to the atomic plane [15]. The Gilbert damping $\alpha$ of 1QL Fe3GaTe2, calculated using Eq. (1), is plotted in Fig. 1(b) as a function of electronic scattering rate $\gamma$. With out-of-plane magnetization ($\mathbf{m}\|\langle 0001\rangle$), $\alpha$ increases monotonically with increasing $\gamma$. Detailed analysis shows that the intraband transitions are completely absent, leading to vanishing conductivity-like behavior. In contrast, for in-plane magnetization ($\mathbf{m}\|\langle 2\bar{1}\bar{1}0\rangle$), $\alpha$ exhibits a nonmonotonic dependence on $\gamma$, with both conductivity-like and resistivity-like terms contributing. Notably, the significant variation in conductivity-like damping at low scattering rates suggests that a substantial anisotropy in Gilbert damping can be achieved by rotating the magnetization.
@@
The absence of conductivity-like damping with out-of-plane magnetization can be explained through a symmetry analysis of the electronic states. We define $\mathcal{M}_{z}$ as the mirror operator [43] about the central plane of the 1QL Fe3GaTe2. For $\mathbf{m}\|\langle 0001\rangle$, the system is mirror-symmetric, and thus $\mathcal{M}_{z}$ commutes with the total Hamiltonian, as well as the SOC Hamiltonian $H_{\textrm{soc}}$ [43]. Consequently, each Bloch state $\Psi_{n\mathbf{k}}$ of the system is an eigenfunction of $\mathcal{M}_{z}$ with eigenvalues $a_{n\mathbf{k}}=-i$ (symmetric) or $+i$ (antisymmetric). These mirror symmetry properties of Bloch states for 1QL Fe3GaTe2 are explicitly calculated using the IRVSP code [44], as shown in Fig. 2(a), where red and blue lines represent symmetric and antisymmetric energy bands with out-of-plane magnetization. Since $\mathcal{M}_{z}$ anti-commutes with $\sigma^{-}$, we have $\{\mathcal{M}_{z},\,[\sigma^{-},H_{\rm soc}]\}=0$. Considering the mirror symmetry properties of Bloch states, we find
$\Gamma^{-}_{mn}(\mathbf{k})=-a^{*}_{mk}a_{nk}\Gamma^{-}_{mn}(\mathbf{k}).$ (2)
@@
For intraband transitions ($m=n$), $|a_{n\mathbf{k}}|^{2}=1$ always holds, leading to $\Gamma^{-}_{nn}(\mathbf{k})=-\Gamma^{-}_{nn}(\mathbf{k})$. This results in $\Gamma^{-}_{nn}(\mathbf{k})=0$ indicating that the intraband transitions are completely forbidden, regardless of the symmetry of the Bloch states. Consequently, the conductivity-like damping vanishes.
@@
Figure 2: (a) Calculated band structure of 1QL Fe3GaTe2 with out-of-plane magnetization. The red and blue lines denote mirror-symmetric ($a_{n\mathbf{k}}=-i$) and antisymmetric ($a_{n\mathbf{k}}=+i$) bands, respectively. Bands calculated with in-plane magnetization are presented in gray. (b) Schematic illustration of energy dissipation during magnetization dynamics in a mirror-symmetric system. (c) Enlarged view of a typical symmetry-protected band crossing, as indicated by the rectangle in panel (a).
@@
Alternatively, the vanishing conductivity-like damping can be intuitively understood by examining the energy dissipation during magnetization dynamics within Kamberský’s breathing Fermi surface model [37, 40]. When the magnetization is out-of-plane at equilibrium, the transverse (in-plane) components of the spin angular momentum $\Delta\mathbf{S}$ relax, as illustrated in Fig. 2(b). The energy of every Bloch state changes with the spin variation $\Delta\mathbf{S}$ through the SOC, i.e., $\Delta\epsilon_{n\mathbf{k}}=\langle\Delta H_{\rm soc}\rangle_{n\mathbf{k}}\propto\langle\mathbf{L}\rangle_{n\mathbf{k}}\cdot\Delta\mathbf{S}$. Due to the mirror symmetry, the orbital angular momentum must have the same out-of-plane component but opposite in-plane components on either side of the mirror plane, namely $L_{z}(z)=L_{z}(-z)$ and $L_{x,y}(z)=-L_{x,y}(-z)$. Therefore, after integrating over the entire space, the band energy remains unchanged with relaxation of the in-plane spin angular momentum, $\Delta\epsilon_{n\mathbf{k}}=0$. This implies that the breathing Fermi surface during the magnetization dynamics does not excite or dissipate the band energy. Consequently, the conductivity-like damping is absent.
@@
To further confirm that the vanishing conductivity-like damping is indeed a result of symmetry, we intentionally break the mirror symmetry in two distinct ways. First, we gradually rotate the magnetization of 1QL Fe3GaTe2 away from the surface normal. This rotation means the Bloch states are no longer eigenstates of $\mathcal{M}_{z}$. The calculated energy bands for in-plane magnetization are plotted with gray lines in Fig. 2(a), where the hybridization of the red and blue bands forms anticrossing gaps. A typical example of the energy bands is shown in the zoomed-in plot in Fig. 2(c). Such hybridization of energy bands facilitates intraband transitions, leading to the emergence of the conductivity-like damping, corresponding to the black squares in Fig. 1(b). We select two scattering rates, $\gamma=3$ meV and 20 meV, as marked by the vertical lines in Fig. 1(b), and calculate damping as a function of the tilting angle $\theta$. As shown in Fig. 3(a), the calculated $\alpha$ increases monotonically with $\theta$ and tends to saturate at $\theta>60^{\circ}$. The saturation suggests that the energy bands hybridization is strong enough to no longer be treated as a perturbation. The enhanced damping is dominated by the intraband transitions (the dotted lines) for both $\gamma$ values, while the interband contributions (the dashed lines) remain nearly unchanged.
The second approach involves varying the number of stacking layers of Fe3GaTe2. By analyzing the lattice structures of 1QL, 2QL and 3QL Fe3GaTe2, we find that mirror symmetry is preserved for odd-number stacking layers but broken in the 2QL system. The calculated damping of Fe3GaTe2 with different numbers of stacking layers is shown in Fig. 3(b), plotted as a function of the simultaneously calculated electrical resistivity $\rho$. The resistivity is calculated using the Boltzmann equation within the relaxation time approximation, where $\rho$ is proportional to the scattering rate $\gamma=\hbar/2\tau$. The calculated $\alpha$ for 1QL and 3QL both monotonically increase with $\rho$, indicating the absence of conductivity-like damping, as expected due to mirror symmetry. In contrast, $\alpha$ of 2QL Fe3GaTe2 exhibits a nonmonotonic dependence on $\rho$ with noticeable conductivity-like contributions at low $\rho$. The calculated $\alpha$ for bulk Fe3GaTe2, shown as the black squares in Fig. 3(b), contains a much larger conductivity-like component. This is because the mirror symmetry is well defined only for the Bloch states with $k_{z}=0$ and those $k_{z}$ located at the boundary plane of the first Brillouin zone [43]. Other electronic states between these special planes are not eigenstates of the mirror operator. Compared with the bulk case, the significantly reduced damping in Fe3GaTe2 thin films at low $\rho$ manifests the strong suppression of the intraband transitions and thus the conductivity-like behavior.
@@
Figure 3: (a) Gilbert damping for 1QL Fe3GaTe2 as a function of the tilting angle $\theta$ of the magnetization away from the surface normal, for two scattering rates $\gamma=3$ meV and 20 meV. The intraband and interband contributions to the damping are represented by dotted and dashed curves, respectively. (b) Gilbert damping for Fe3GaTe2 with varying stacking layers plotted as a function of electric resistivity.
@@
Interband transitions.—Although intraband transitions are prohibited by mirror symmetry, Eq. (2) indicates that interband transitions are allowed between Bloch states with contrasting symmetry characteristics, i.e., $a_{m\mathbf{k}}\neq a_{n\mathbf{k}}$. In fact, the mirror symmetry ensures the presence of nodal lines in the Brillouin zone, which are formed by the intersections of mirror-symmetric and mirror-antisymmetric bands [45, 46]. According to Eq. (1), the interband contribution to damping depends on the overlap of the broadened energy bands at the Fermi energy $E_{F}$. As schematically illustrated in Fig. 4(a), the interband contribution from the two intersecting bands is maximized when the intersection point aligns with the Fermi level, i.e., $E_{F}=E_{3}$. Consequently, pronounced interband transitions are anticipated near band crossings or nodal lines close to $E_{F}$. In the case of 1QL Fe3GaTe2, the nodal lines in the 2D Brillouin zone are predominantly located several hundred meV above $E_{F}$, as shown in Fig. 4(b). This accounts for the very low interband damping of 1QL Fe3GaTe2 with out-of-plane magnetization, particularly at low $\gamma$.
@@
Figure 4: (a) Schematic representation of the overlap between two broadened energy bands near their crossing point, with shadows indicating the scattering rate-induced broadening. Thick black lines show the extent of overlap at various isoenergetic surfaces. (b) Calculated nodal lines for 1QL Fe3GaTe2 with out-of-plane magnetization in the 2D Brillouin zone. (c) Gilbert damping $\alpha$ as a function of energy at two scattering rates. (d) Gilbert damping $\alpha$ as a function of scattering rate at various energy levels.

### v3

# Symmetry-forbidden intraband transitions leading to ultralow Gilbert damping in van der Waals ferromagnets

@@ Abstract
Based upon first-principles calculations, we report ultralow Gilbert damping in two-dimensional (2D) van der Waals (vdW) ferromagnets. The low damping occurs at weak scattering because mirror symmetry prohibits intraband transitions. The monotonic dependence on the electronic scattering rate suggests the absent lower limit, in contrast to conventional ferromagnetic materials. Breaking mirror symmetry through magnetization rotation, layer stacking, or structural phase transition significantly increases damping by enabling intraband transitions. Topological nodal lines, also protected by mirror symmetry, contribute substantially to interband-transition-mediated damping, which can be tuned by adjusting the Fermi level. Our findings elucidate the unique characteristics of Gilbert damping in 2D vdW ferromagnets, providing valuable insights for designing low-dimensional spintronic devices with high energy efficiency.
@@
Introduction.—The discovery of 2D vdW ferromagnetic (FM) materials, exemplified by CrI3 [1], Cr2Ge2Te6 [2], and Fe3GeTe2 [3, 4], has not only dramatically broadened the horizon of FM materials but also ignited promising prospects for low-dimensional spintronics devices [5, 6]. This promise stems partly from the fact that magnetism in 2D vdW FM metals can be controlled through unique approaches like gating and strain [7, 8]. As a typical 2D FM metal, Fe3GeTe2 has attracted considerable attention for its spin-orbit torque-driven magnetization switching [9, 10, 11], its unconventional tunability in magnetic tunnel junctions [12, 13] and its potential in novel electronic applications [14]. A very similar vdW FM metal Fe3GaTe2 [15] has been discovered with an almost identical lattice structure to Fe3GeTe2, yet it boasts a Curie temperature above room temperature, rendering it an ideal candidate for integration into magnetic heterostructures and devices that operate under ambient conditions [16, 17, 18, 19].
@@
The quest for designing high-performance, low-consumption spintronics devices based on 2D vdW FM materials requires a comprehensive understanding of the dynamical properties of 2D magnetism [20], particularly the effects of reduced dimensionality. Gilbert damping, a crucial parameter in magnetization dynamics [21], determines the critical current density for current-induced magnetization switching [22], the switching time scale [23] and the velocity of magnetic texture driven by magnetic fields and/or electric currents [24, 25]. More importantly, low Gilbert damping is essential for low-consumption in memory and logic devices. Ultralow damping has been observed in Fe-based alloys [26, 27, 28, 29] and half metals [30, 31]. In FM transition metals and alloys, such as Fe, Co, Ni [32, 33, 34] and FePd [35, 36], the experimentally measured damping $\alpha$ exhibits a non-monotonic temperature dependence: $\alpha$ increases with decreasing temperature (conductivity-like) at low-temperatures and with increasing temperature (resistivity-like) at high-temperatures, limiting damping above its lower bound. However, in the recently discovered 2D FM metals, Fe3GeTe2 and Fe3GaTe2, a monotonic temperature dependence of damping has been observed in both experimental [37] and theoretical studies [38]. Furthermore, significant damping anisotropy has been identified in these 2D FM metals [38, 39], suggesting an atypical magnetization relaxation likely due to their low-dimensional lattice structures.
@@
In this Letter, we present a theoretical calculation of the Gilbert damping in 2D vdW FM metals, using Fe3GaTe2 as a typical example. We demonstrate that mirror symmetry in 2D vdW systems prohibits intraband transitions, leading to the absence of conductivity-like damping. Consequently, the damping exhibits a monotonic temperature dependence without the lower limit. An ultralow $\alpha<10^{-3}$ is achieved in the single layer at low temperatures. Furthermore, symmetry-protected nodal lines near the Fermi level substantially enhance interband transitions, thereby increasing the resistivity-like damping. Our findings predict that the damping can be reduced to arbitrarily low values by decreasing the electronic scattering rate. These unique features are generally applicable to other 2D vdW materials with mirror symmetry, such as Fe3GeTe2 and 2H-FeTe2.
@@
Theoretical methods.—We utilize the torque-correlation model [40, 41, 42] to calculate the Gilbert damping of 2D vdW FM metals based upon first-principles electronic structure. This model captures the dynamical dissipation of magnetization due to spin-orbit coupling (SOC), corresponding to the dominant contribution to the intrinsic Gilbert damping $\alpha$ [43]. Explicitly, $\alpha$ is given by
$\alpha=\frac{\pi g\mu_{\textrm{B}}}{M_{s}}\sum_{n,m}\int\frac{d^{3}k}{(2\pi)^{3}}|\Gamma^{-}_{mn}(\mathbf{k})|^{2}\int d\epsilon\eta(\epsilon)A_{n\mathbf{k}}(\epsilon)A_{m\mathbf{k}}(\epsilon).$ (1)
@@
Here $g$, $\mu_{\textrm{B}}$ and $M_{s}$ are the Landé factor, the Bohr magneton and the saturation magnetization, respectively. While indices $n$ and $m$ run over all energy bands, the energy derivative of the Fermi-Dirac distribution $\eta(\epsilon)\equiv-\partial f/\partial\epsilon$ ensures that the damping is predominantly influenced by electronic states near the Fermi surface. The matrix element $\Gamma^{-}_{mn}(\mathbf{k})=\langle\Psi_{m\mathbf{k}}|[\sigma^{-},H_{\textrm{soc}}]|\Psi_{n\mathbf{k}}\rangle$ characterizes the torque on electron spin stemming from SOC. The spectral function $A_{n\mathbf{k}}(\epsilon)$ is modeled as a Lorentzian centered at the band energy $\epsilon_{n\mathbf{k}}$, with its width determined by the electronic scattering rate $\gamma$ [41, 42]. Prior studies have shown that intraband ($n=m$) transitions yield a conductivity-like damping component, which is proportional to the electronic relaxation time, $\tau=\hbar/2\gamma$ [40, 41], while interband ($n\neq m$) transitions contribute to a resistivity-like damping component that scales monotonically with increasing $\gamma$ [44]. To explicitly evaluate these intraband and interband transitions in Eq. (1), we have developed a Wannier interpolation method to perform the integral of $|\Gamma^{-}_{mn}(\mathbf{k})|^{2}$ based on the first-principles electronic structures. The self-consistent band structures are calculated using Quantum ESPRESSO [45], and the Wannier functions are constructed using WANNIER90 [46]. Here, we retain only the SOC-induced Gilbert damping that governs uniform magnetization precession, as measured by ferromagnetic resonance (FMR) experiments. The Stoner damping [47, 48, 49] mediated by exchange interaction that contributes to nonuniform magnetization dynamics, and the two-magnon scattering [50, 51] that can be technically excluded from FMR measurements, are beyond the scope of the present work.
Figure 1: (a) Atomic structure of 1QL Fe3GaTe2. The magnetic moments on the Fe atoms are indicated by black arrows. The blue plane in the middle of the layer highlights the mirror symmetry of the lattice structure. (b) Gilbert damping as a function of the electronic scattering rate for 1QL Fe3GaTe2 with the out-of-plane (red solid circles) and in-plane magnetization (black solid squares). The dotted and dashed lines represent the contributions from intraband and interband transitions, respectively. The orange and green vertical lines denote two characteristic scattering rates, $\gamma=3$ meV and 20 meV, which are used to calculate the damping in detail.
@@
Intraband transitions.— We choose Fe3GaTe2 as a typical example to illustrate the unique characteristics of Gilbert damping in 2D vdW ferromagnets. The lattice structure of a single layer Fe3GaTe2, referred to as one quintuple-layer (1QL), is depicted in Fig. 1(a). This lattice, belonging to the D3h point group, maintains mirror symmetry about the blue plane only when the magnetization is perpendicular to the atomic plane [15]. The Gilbert damping $\alpha$ of 1QL Fe3GaTe2, calculated using Eq. (1), is plotted in Fig. 1(b) as a function of electronic scattering rate $\gamma$. With out-of-plane magnetization ($\mathbf{m}\|\langle 0001\rangle$), $\alpha$ increases monotonically with increasing $\gamma$. Detailed analysis shows that the intraband transitions are completely absent, leading to vanishing conductivity-like behavior. As a result, $\alpha$ is lower than $10^{-3}$ at small scattering rate $\gamma<10$ meV. In contrast, for in-plane magnetization ($\mathbf{m}\|\langle 2\bar{1}\bar{1}0\rangle$), $\alpha$ exhibits a nonmonotonic dependence on $\gamma$, with both conductivity-like and resistivity-like terms contributing.
@@
The ultralow damping due to the absence of conductivity-like damping with out-of-plane magnetization can be explained through a symmetry analysis of the electronic states. We define $\mathcal{M}_{z}$ as the mirror operator [52] about the central plane of the 1QL Fe3GaTe2. For $\mathbf{m}\|\langle 0001\rangle$, the system is mirror-symmetric, and thus $\mathcal{M}_{z}$ commutes with the total Hamiltonian, as well as the SOC Hamiltonian $H_{\textrm{soc}}$ [52]. Consequently, each Bloch state $\Psi_{n\mathbf{k}}$ of the system is an eigenfunction of $\mathcal{M}_{z}$ with eigenvalues $a_{n\mathbf{k}}=-i$ (symmetric) or +i+i (antisymmetric). These mirror symmetry properties of Bloch states for 1QL Fe3GaTe2 are explicitly calculated using the IRVSP code [53], as shown in Fig. 2(a), where red and blue lines represent symmetric and antisymmetric energy bands with out-of-plane magnetization. Since $\mathcal{M}_{z}$ anti-commutes with $\sigma^{-}$, we have $\{\mathcal{M}_{z},\,[\sigma^{-},H_{\rm soc}]\}=0$. Considering the mirror symmetry properties of Bloch states, we find
$\Gamma^{-}_{mn}(\mathbf{k})=-a^{*}_{m\mathbf{k}}a_{n\mathbf{k}}\Gamma^{-}_{mn}(\mathbf{k}).$ (2)
@@
For intraband transitions ($m=n$), $|a_{n\mathbf{k}}|^{2}=1$ always holds, leading to $\Gamma^{-}_{nn}(\mathbf{k})=-\Gamma^{-}_{nn}(\mathbf{k})$. Thus we have $\Gamma^{-}_{nn}(\mathbf{k})=0$ indicating that the intraband transitions are completely forbidden, regardless of the symmetry of the Bloch states. Consequently, the conductivity-like damping vanishes and $\alpha$ increases monotonically with $\gamma$.
@@
Figure 2: (a) Calculated band structure of 1QL Fe3GaTe2 with out-of-plane magnetization. The red and blue lines denote mirror-symmetric ($a_{n\mathbf{k}}=-i$) and antisymmetric ($a_{n\mathbf{k}}=+i$) bands, respectively. Bands calculated with in-plane magnetization are presented in gray. (b) Schematic illustration of energy dissipation during magnetization dynamics in a mirror-symmetric system. (c) Enlarged view of a typical symmetry-protected band crossing, as indicated by the rectangle in (a).
@@
Alternatively, the symmetry-forbidden intraband transitions can be intuitively understood by examining the energy dissipation during magnetization dynamics within Kamberský’s breathing Fermi surface model [40, 44]. When the magnetization is out-of-plane at equilibrium, the transverse (in-plane) components of the spin angular momentum $\Delta\mathbf{S}$ relax, as illustrated in Fig. 2(b). The energy changes with the spin variation $\Delta\mathbf{S}$ through SOC, i.e., $\Delta E=\Delta H_{\rm soc}\propto\mathbf{L}\cdot\Delta\mathbf{S}$. Due to the mirror symmetry, the orbital angular momentum $\mathbf{L}$ must have the same out-of-plane component but opposite in-plane components on either side of the mirror plane, namely $L_{z}(z)=L_{z}(-z)$ and $L_{x,y}(z)=-L_{x,y}(-z)$. Therefore, after integrating over the entire space, the band energy remains unchanged with the in-plane $\Delta\mathbf{S}$, i.e., $\Delta E=0$. This implies that the breathing Fermi surface during the magnetization dynamics does not excite the system out of equilibrium, and the intraband transitions are unnecessary to bring it back to equilibrium.
@@
To further confirm that the absent intraband transitions (conductivity-like damping) is indeed a result of symmetry, we intentionally break the mirror symmetry in different ways. First, we rotate the magnetization of 1QL Fe3GaTe2 away from the surface normal. This rotation means the Bloch states are no longer eigenstates of $\mathcal{M}_{z}$. The calculated energy bands for in-plane magnetization are plotted with gray lines in Fig. 2(a), where the hybridization of the red and blue bands forms anticrossing gaps with a typical example magnified in Fig. 2(c). Such band hybridization facilitates intraband transitions, leading to the emergence of the conductivity-like damping due to intraband transitions, corresponding to the black squares in Fig. 1(b). We select two scattering rates, $\gamma=3$ meV and 20 meV, as marked by the vertical lines in Fig. 1(b), and calculate damping as a function of the tilting angle $\theta$. As shown in Fig. 3(a), the calculated $\alpha$ increases monotonically with $\theta$ and tends to saturate at $\theta>60^{\circ}$. The saturation suggests that the energy bands hybridization is strong enough to no longer be treated as a perturbation. The enhanced damping is dominated by the intraband transitions (the dotted lines) for both $\gamma$ values, while the interband contributions (the dashed lines) remain nearly unchanged. We conducted an analysis on Fe3GeTe2, whose lattice structure is identical to Fe3GaTe2, except for the substitution of Ga atoms with Ge. For 1 QL Fe3GeTe2 with perpendicular magnetization, conductivity-like damping vanishes, resulting in a monotonic increase of the calculated $\alpha$ with increasing $\gamma$. Rotating the magnetization leads to recovery of intraband contributions, thus yielding non-monotonic behavior, which is qualitatively the same as Fe3GaTe2; see Sec. 1 in Supplemental Material [54].
@@
Figure 3: (a) Gilbert damping for 1QL Fe3GaTe2 as a function of the tilting angle $\theta$ of the magnetization away from the surface normal, for two scattering rates $\gamma=3$ meV and 20 meV. The intraband and interband contributions to the damping are represented by dotted and dashed curves, respectively. (b) Gilbert damping for Fe3GaTe2 with varying stacking layers plotted as a function of electrical resistivity.
The second approach involves varying the number of stacking layers of Fe3GaTe2. By analyzing the lattice structures of 1QL, 2QL and 3QL Fe3GaTe2, we find that mirror symmetry is preserved for odd-number stacking layers but broken in the 2QL system. The calculated damping of Fe3GaTe2 with different numbers of stacking layers is shown in Fig. 3(b), plotted as a function of the simultaneously calculated electrical resistivity $\rho$. The resistivity is calculated using the Boltzmann equation within the relaxation time approximation, where $\rho$ is proportional to the scattering rate $\gamma=\hbar/2\tau$. The calculated $\alpha$ for 1QL and 3QL both monotonically increase with $\rho$, indicating the forbidden intraband transitions, as expected due to mirror symmetry. In contrast, $\alpha$ of 2QL Fe3GaTe2 exhibits a nonmonotonic dependence on $\rho$ with noticeable conductivity-like damping arising from intraband transitions at low $\rho$. The calculated $\alpha$ for bulk Fe3GaTe2, shown as the black squares in Fig. 3(b), contains a much larger conductivity-like component. This is because the mirror symmetry is well defined only for the Bloch states with $k_{z}=0$ and those $k_{z}$ located at the boundary plane of the first Brillouin zone [52]. Other electronic states between these special planes are not eigenstates of the mirror operator and intraband transitions are not prohibited by mirror symmetry. Compared with the bulk case, the significantly reduced damping in Fe3GaTe2 thin films at low $\rho$ manifests the strong suppression of the intraband transitions and thus the conductivity-like behavior.
We further performed calculations for 2D monolayer FeTe2 in two distinct structural phases: the trigonal prismatic 2H phase and the octahedral 1T phase. The mirror-symmetric 2H-FeTe2 with out-of-plane magnetization exhibits monotonic damping. In contrast, the damping in 1T-FeTe2, which lacks mirror symmetry, behaves non-monotonically and is independent of magnetization orientation; see Sec. 2 in Supplemental Material [54]. These additional calculations indicate that the symmetry-forbidden intraband transition is a general feature of any mirror-symmetric 2D vdW ferromagnets.
@@
Interband transitions.—Although intraband transitions are prohibited by mirror symmetry, Eq. (2) indicates that interband transitions are allowed between Bloch states with contrasting symmetry characteristics, i.e., $a_{m\mathbf{k}}\neq a_{n\mathbf{k}}$. In fact, the mirror symmetry ensures the presence of nodal lines in the Brillouin zone, which are formed by the intersections of mirror-symmetric and mirror-antisymmetric bands [55, 56]. According to Eq. (1), the interband contribution to damping depends on the overlap of the broadened energy bands at the Fermi energy $E_{F}$. As schematically illustrated in Fig. 4(a), the interband contribution from the two intersecting bands is maximized when the intersection point aligns with the Fermi level, i.e., $E_{F}=E_{3}$. Consequently, pronounced interband transitions are anticipated near band crossings or nodal lines close to $E_{F}$. In the case of 1QL Fe3GaTe2, the nodal lines in the 2D Brillouin zone are located several hundred meV above $E_{F}$, as shown in Fig. 4(b). This accounts for the very low interband damping of 1QL Fe3GaTe2 with out-of-plane magnetization, particularly at low $\gamma$.
@@
Figure 4: (a) Schematic representation of band overlap near their crossing point, with shadows indicating the broadening. Thick black lines show the extent of overlap at various isoenergetic surfaces. (b) Calculated nodal lines for 1QL Fe3GaTe2 with out-of-plane magnetization in the 2D Brillouin zone. (c) Calculated $\alpha$ as a function of energy at two scattering rates. (d) Calculated $\alpha$ as a function of scattering rate at various energies.